OpenStudy (lilsis76):

Help me understand. Determine the equilibrium constant for the following reaction at 398 K. 3Cu(s) + 2MnO4- (aq) + 8H+(aq) --> 3Cu2+(aq) + 2MnO2(s) + 4H2O (l)

OpenStudy (aaronq):

hm have you used the nernst equation?

OpenStudy (lilsis76):

yes we have before but the instructor, she goes thru it so fast in her slides that I can hardly keep up an we have an exam on tuesday @aaronq

OpenStudy (lilsis76):

@aaronq its that E cell = ECell standard potential -.0592V/n logQ

OpenStudy (aaronq):

at equilibrium \((i.e.~E_{cell}=E^o_{cell})\), so you can solve for K. \(0=E^o_{cell}-\dfrac{RT}{nF}lnK\rightarrow E^o_{cell}=\dfrac{RT}{nF}lnK\rightarrow lnK=\dfrac{nFE^0_{cell}}{RT}\)

OpenStudy (lilsis76):

wha?! haha but do i have k?

OpenStudy (aaronq):

K is the equilibrium constant (what you're looking for).

OpenStudy (aaronq):

remember that the reduced equation you wrote can only be used at 298 K, 1M concentration and at 1 atm (i.e. standard conditions) which is not the case here (398 K).

OpenStudy (lilsis76):

oh...okay haha so what do I do?

OpenStudy (lilsis76):

@aaronq im looking at my note paper with equtions there is a Ecell=.0592V/n * log K and there is a Ecell = Ecell- .0592V/n *logQ

OpenStudy (aaronq):

use the original form: \(E_{cell}=E^0_{cell}-\dfrac{RT}{nF}lnK\) , which solves to what i wrote above

OpenStudy (lilsis76):

okay. so then... here hold on .. the nF = hold on I have to do that reduction oxidation thingy right? and for that oxidation would be anode so negative, an cathode is reduction that is positive right? Im just making sure I have this part right first. @aaronq

OpenStudy (lilsis76):

I also found this chart just in case http://bilbo.chm.uri.edu/CHM112/tables/redpottable.htm

OpenStudy (lilsis76):

UGH!!! IM SO LOST! okay this is what I understand from your formula: nF= the number of mol electrons R= the constant, i think its 8.314 J/mol K T= is the temperature=398K @aaronq Is all of this right so far? okay I just dont remember how to find the \[E^0 cell\]

OpenStudy (aaronq):

n = # of electrons exchanged in redox reaction: 3Cu(s) + 2MnO4- (aq) + 8H+(aq) --> 3Cu2+(aq) + 2MnO2(s) + 4H2O (l) 6 electrons

OpenStudy (aaronq):

F = faradays constant

OpenStudy (lilsis76):

okay F = Faradays constant is 96485 C/mol e-

OpenStudy (aaronq):

yep, thats right. now you just need to find the potential of the cell \(E^0_{cell}=E^{red}_{cathode}-E^{red}_{anode}\)

OpenStudy (aaronq):

sometimes it's also written as \(E^0_{cell}=E^{red}_{cathode}+E^{oxd}_{anode}\)

OpenStudy (lilsis76):

oh....okay wait haha i always thought u would subtract it, but I guess I keep getting it wrong, so basically it becomes an addition

OpenStudy (aaronq):

you subtract it if both of your values are reduction potentials. you add them if one is oxidation (at anode) and one is reduction potential (at cathode).

OpenStudy (lilsis76):

okay so @aaronq the anode would be the positive charge right? or am i mixed up. I remember An-Ox for Anode=Oxidation= - and Red-Cat for Cathode=Reduction=+

OpenStudy (aaronq):

\(\huge \color{red}OI\color{red}L ~~\color{blue}RI\color{blue}G\) Oxidation loss of electrons (at anode) Reduction gain of electrons (at cathode)

OpenStudy (lilsis76):

okay let me see if I have this right then

OpenStudy (lilsis76):

SORRY @aaronq INTERNET CONNECTION IS being lame

OpenStudy (aaronq):

no worries !

OpenStudy (lilsis76):

okay the Cathode, or reduction reaction is going to be the 2MnO4

OpenStudy (aaronq):

yes

OpenStudy (lilsis76):

right? cuz it goes from a 2MnO4- to a 2MnO2; it ganed the e- to make it a positive and added Oxygen :D

OpenStudy (lilsis76):

ok, so then now that i know that

OpenStudy (lilsis76):

okay E cell.er....um..... okay hold on

OpenStudy (aaronq):

sorry, i read that as "now what" lol

OpenStudy (lilsis76):

Would I use....half reactions for the Ecell(V) ?

OpenStudy (aaronq):

yes

OpenStudy (aaronq):

so find these reactions in a reduction potential chart and set them up

OpenStudy (lilsis76):

http://chemed.chem.purdue.edu/genchem/topicreview/bp/ch20/electro.php Im using this website

OpenStudy (lilsis76):

Cu2+ + e- <-----> Cu+ 0.158 so I would flip it right? and MnO4- + 2 H2O + 3 e- <-----> MnO2 + 4 OH- 0.588

OpenStudy (lilsis76):

@aaronq

OpenStudy (aaronq):

you're using this one Cu2+ + 2 e- <-----> Cu 0.3402

OpenStudy (aaronq):

but yeah flip that one

OpenStudy (lilsis76):

wait haha I just forgot, but why do I flip it?

OpenStudy (lilsis76):

and why do I use that one an not the others? @aaronq

OpenStudy (lilsis76):

@aaronq I also cant figure out which MnO4 to use: MnO4- + 2 H2O + 3 e- <-----> MnO2 + 4 OH- 0.588 or MnO2 + 4 H+ + 2 e- <-----> Mn2+ + 2 H2O 1.208 or MnO4- + 8 H+ + 5 e- <-----> Mn2+ + 4 H2O 1.491

OpenStudy (lilsis76):

well in the main equation they have an 8H+ --> 4H+ would I use the bottom one that gives me 1.491? @aaronq

OpenStudy (aaronq):

you flip it if you're adding the values together, because as it's written it is a reduction equation and you're looking for the oxidation.

OpenStudy (aaronq):

i would use the top one from those 3

OpenStudy (lilsis76):

ohh...okay haha @aaronq thank you so then is this right, okay dang hold on gotta do it again but not use the bottom number

OpenStudy (aaronq):

no i mean the bottom one. use the bottom one of the 3

OpenStudy (lilsis76):

OKAY, oh........haha okay let me do it again hold on haha

OpenStudy (lilsis76):

okay @aaronq I did the Ecell= C+A 1.491+-.34 = 1.151v

OpenStudy (lilsis76):

right?

OpenStudy (lilsis76):

right?

OpenStudy (aaronq):

yep thats right

OpenStudy (lilsis76):

@aaronq sorry I keep loosing connection, is ur computer the same way?

OpenStudy (aaronq):

no, i was just away from the computer. so, \(\huge K=e^{\small(\dfrac{nFE^o_{cell}}{RT})}\)

OpenStudy (lilsis76):

OH! ahah u rearranged that one equation huh? haha

OpenStudy (aaronq):

haha yeah, you should try it yourself on paper to make sure you can arrive at the same result

OpenStudy (lilsis76):

@aaronq Im going to need your help big time right now because. I dont remember the n, you said the n was the number of elections exchanged in redox reaction. u said that was 6 because there was 3 Cu and 2MnO4's so you times them and got 6 right?

OpenStudy (aaronq):

yes

OpenStudy (lilsis76):

okay then so.. @aaronq

OpenStudy (lilsis76):

e=1.151v

OpenStudy (lilsis76):

\[1.151v (\frac{ 8.314*398K }{ 6*96485 })lnK\]

OpenStudy (lilsis76):

oops I think I did it wrong @aaronq

OpenStudy (lilsis76):

\[k=1.151v(\frac{ 6*96485*e^0cell }{ 8.314*398K })\] Right ????@aaronq

OpenStudy (lilsis76):

@aaronq is the answer....Well I dont know what to put for the e^0 Cell :/

OpenStudy (lilsis76):

this dumb connection! @aaronq

OpenStudy (aaronq):

so you had to end up with: \(\sf K=e^\dfrac{6*(9.65*10^4 C/mol)*1.151V}{(8.314 J/mol*K)*(398K)}\)

OpenStudy (lilsis76):

im back @aaronq

OpenStudy (lilsis76):

yes I did @aaronq but what do i put for the e??

OpenStudy (lilsis76):

@aaronq I get K=e 201.4

OpenStudy (lilsis76):

do i use that 10^?? to get.... WOW, haha I got an over flow error

OpenStudy (aaronq):

"e" is euler's number. it's transcendental but you can write it as 2.718 so it's \(2.718^{2.93*10^{87}}\) i'm getting \(\infty\), which just means that the number is too big for my calculator lol

OpenStudy (lilsis76):

okay @aaronq i got a big number haha. I dont even think we were doing that part right. Here is what I am given on an equation sheet of mine

OpenStudy (lilsis76):

I will be back online in the morning.

OpenStudy (aaronq):

i think it was late and i was doing something wrong. now ii'm getting \(e^{201.4}=2.93*10^{87}\) which seems right.

OpenStudy (aaronq):

ps i'm 100% sure were using the right equation.

OpenStudy (lilsis76):

haha @aaronq okay Im back! had to go out to store.

OpenStudy (lilsis76):

@aaronq but where okay hold on so this was the equation we were working on: well this part: \[K=e(\frac{ 6*96485*1.151v }{ 8.314*398K })\] giving us: \[K=e(\frac{ 666325.4 }{ 8.314*398K }) \] \[K=e201.4\] @aaronq NOW, do we raise both to the e^, but how does it get rid of the K? I understand how u got\[K=e^{201.4} = 2.93x10^{87}\]

OpenStudy (aaronq):

when you have: \(lnK=\dfrac{nFE}{RT}\) you have to raise both sides to the \(e^x\) to get rid of ln. so you end up with \(\Huge K=e^{\small \frac{nFE}{RT}}\) e^(nFE/RT)

OpenStudy (lilsis76):

oh. okay that makes sense now lol sorry

OpenStudy (aaronq):

the formatting its kinda weird because there are so many symbols in the exponent, which makes like it look like a multiplication, but it's not.

OpenStudy (aaronq):

no problem ! hope everything is good now

OpenStudy (lilsis76):

okay so now that we have the K=2.93x10^87 that will be our volts, right?

OpenStudy (aaronq):

no that's the equilibrium constant (usually unitless)

OpenStudy (lilsis76):

really? so we just solved it? I thought there would need another equation

OpenStudy (aaronq):

hah yep thats it

OpenStudy (lilsis76):

Dang....Longest problem ever

OpenStudy (lilsis76):

thank you so much. I have other problems too that I will post on here in a little bit.

OpenStudy (aaronq):

no problem !